OpenStudy (anonymous):

simplify radical 48??

OpenStudy (hawkfalcon):

What fits into it that can be rooted?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

use \[\Large \sqrt{16*3}=\sqrt{16} \sqrt{3}\] i hope you can do this now.

OpenStudy (anonymous):

\[\sqrt{48}=\sqrt{4^{2} \cdot 3}=4 \cdot \sqrt{3}\]

OpenStudy (hawkfalcon):

It is exactly like the other one... Find a rootable number that fits evenly within the radical and divide it by that.

OpenStudy (anonymous):

4 radical 3

OpenStudy (anonymous):

yes.

OpenStudy (hawkfalcon):

yes.